OpenStudy (anonymous):

find the area of a shaded circle if the radius is 3/8ths inches

OpenStudy (anonymous):

Gotta' have the picture of the circle. But, the area of the circle is: pi * r^2 = pi * 9 / 64 inches.

OpenStudy (anonymous):

9 * pi / 64 inches

OpenStudy (anonymous):

\[A = \pi r^2 = \pi(3/8)^2 = 9\pi/64 \]

OpenStudy (anonymous):

so 9*3.14 divided by 64?